# Multiple exponential sums and their applications to quadratic congruences

### v3

@@ 3. Proof of Proposition 1
Next, we consider the case when $m-r$ is odd, i.e. $2|(m-r-1)$. The calculations are similar as above, and writing ${\bf{x}}=p^{(m-r+1)/2}{\bf{z}}+{\bf{y}}$ with ${\bf{z}}\in\left(\mathbb{Z}p/p^{(m+r-1)/2}\mathbb{Z}_{p}\right)^{n}$ and ${\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}$, we now obtain the following chain of equations.
$S_{\bf{\alpha}}(f,p^{m})$ $=\sum_{\begin{subarray}{c}{\bf{x}}\in(\mathbb{Z}p/p^{m}\mathbb{Z}_{p})^{n}\\ {\bf{x}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{x}}))$
$=\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m+r-1)/2}\mathbb{Z}_{p})^{n}}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({p^{(m-r+1)/2}\bf{z}}+{\bf y}))$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m+r-1)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(f({\bf{y}})+p^{(m-r+1)/2}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{nr}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m-r-1)/2}\mathbb{Z}_{p})^{n}}e_{p^{(m-r-1)/2}}(p^{-r}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r-1)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})).$
@@ 4. Proof of Theorem 1
$S_{\bf{\alpha}}(f,p^{m})=p^{n(m+r)/2}e_{p^{m}}(f({\bf{\alpha^{*}}})),\quad\mbox{if }m-r\mbox{ is even.}$.
Similarly, if $m-r$ is odd, then $\alpha$ has a unique lifting to a solution ${\bf{\alpha}^{*}}$ of the congruence $p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r+1)/2}$. Now from (1.1), we obtain
$S_{\bf{\alpha}}(f,p^{m})$ $=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p^{m}}(f({\bf{\alpha}^{*}}+P^{(m-r-1)/2}{\bf{u}}))$
$=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p^{m}}(f({\bf{\alpha}^{*}})+P^{(m-r-1)/2}{\bf{u}}\cdot\nabla f({\bf{\alpha}^{*}})+p^{m-r-1}2^{-1}{\bf{u}}^{t}\text{Hesse}f({\bf{\alpha}^{*}}){\bf{u}})$
$=p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(p^{-r}2^{-1}{\bf{u}}^{t}\text{Hesse}f({\bf{\alpha}^{*}}){\bf{u}}).$
Take $A=p^{-r}\text{Hesse}f({\bf{\alpha}^{*}})$, which is a symmetric matrix. Hence there exists an orthogonal matrix, say $P,$ such that,
$PAP^{t}=D=\begin{bmatrix}\lambda_{1}&0&...&0\\ 0&\lambda_{2}&...&0\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&...&\lambda_{n}\end{bmatrix},$
where $\lambda_{1},...,\lambda_{n}$ are the eigenvalues of $A$. Let
@@
where $R_{i}$ are the row vectors of $P$. Now ${\bf u}^{t}A{\bf u}={\bf{u}}^{t}P^{t}DP{\bf{u}}=\lambda_{1}(R_{1}{\bf{u}})^{2}+\lambda_{2}(R_{2}{\bf{u}})^{2}+\cdots+\lambda_{n}(R_{n}{\bf{u}})^{2}$. As $P$ is invertible, the map ${\bf{u}}\mapsto P{\bf u}$ is a bijection on $(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}$. Hence we get
$\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(p^{-r}{\bf{u}}^{t}\text{Hesse}f({\bf{\alpha}^{*}}){\bf{u}})$ $=\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}p/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(\lambda_{1}(R_{1}{\bf{u}})^{2}+\lambda_{2}(R_{2}{\bf{u}})^{2}+\cdots+\lambda_{n}(R_{n}{\bf{u}})^{2})$
@@
$S_{\bf{\alpha}}(f,p^{m})=\left(\frac{\text{det}(p^{-r}\text{Hesse}f({\bf{\alpha}^{*}}))}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))G_{p}(1)^{n},\quad\mbox{if }m-r\mbox{ is odd,}$,
@@ 5.1. Multiple Poisson summation
Setting $x_{i}=y_{i}+k_{i}p^{m}$, where $1\leq y_{i}\leq p^{m}$, we apply Proposition 2 to obtain
@@ 5.2. Evaluation of exponential sums
(5.3) $f_{k_{1},...,k_{n}}(y_{1},y_{2},...,y_{n}):=h\sum_{i,j=1}^{n}a_{ij}y_{i}y_{j}+\sum_{i=1}^{n}k_{i}y_{i}.$
The contributions of $r=m-1$ and $r=m$ to the right-hand side of (5.2) are $O_{\varepsilon}(1)$ if $N\geq p^{m\varepsilon}$ by the rapid decay of $\hat{\Phi}$ as $(k_{1},...,k_{n})\neq(0,...,0)$. We fix $l_{i}=\frac{k_{i}}{p^{r}}$. From (5.3) we calculate
$\nabla f_{k_{1},...,k_{n}}=\left(2h\frac{\partial Q}{\partial y_{1}}+p^{r}l_{1},...,2h\frac{\partial Q}{\partial y_{n}}+p^{r}l_{n}\right).$
If $\text{ord}_{p}(h)>p^{r}$ then $\text{ord}_{p}(\nabla(f))=p^{r}$. So $p^{-r}\nabla(f)\equiv(l_{1},...l_{n})\bmod p,$ where $(l_{1},...,l_{n})\not\equiv(0,...,0)\bmod p$. Hence there is no critical point of $f_{k_{1},...,k_{n}}$. Now using $(i)$ of Theorem 1, we have $E(k_{1},...,k_{n},h;p^{m})=0$. On the other hand if $\text{ord}_{p}(h)<p^{r}$ then $\text{ord}_{p}(\nabla(f))=\text{ord}_{p}(h)$. Hence $p^{\text{ord}_{p}(h)}\nabla(f)\equiv{\boldsymbol{0}}\bmod p,$ which implies $AY\equiv{\boldsymbol{0}}\bmod p$. Since $A$ is non-singular, $Y=(0,...,0)$ is the only critical point in this case. Hence $E(k_{1},...,k_{n},h;p^{m})=0.$ Thus from now onward we consider $\text{ord}_{p}(h)=p^{r}$. Therefore $\text{ord}_{p}(\nabla(f))=p^{r}=(h,p^{m})$ and we write $l=\frac{h}{p^{r}}$. So if $N\geq p^{m\varepsilon}$, we have
(5.4) $U=\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})+O_{\varepsilon}(1).$
@@
(5.5) $\frac{\partial f_{l_{1}p^{r},...,l_{n}p^{r},lp^{r}}}{\partial y_{i}}(y_{1},y_{2},...,y_{n})=p^{r}\left(2l\frac{\partial{Q}}{\partial{y_{i}}}+l_{i}\right),~{}\forall~{}i.$
In the following, we will apply Proposition 1 to evaluate the exponential sums $E(l_{1}p^{r},...,l_{n}p^{r};p^{m})$. Now consider the case when $m-r$ is even. Note that
$p^{-r}\nabla f_{k_{1},...,k_{n}}(y_{1},...,y_{n}))\equiv 0\bmod p^{m-r}$
implies
$2lAY+L\equiv 0\bmod p^{m-r},$
@@
where $Y=(y_{1}~{}y_{2}~{}\cdots~{}y_{n})^{t}$ and $L=(l_{1}~{}l_{2}~{}\cdots~{}l_{n})^{t}$. Now putting the values of $y_{i}$ in (5.3), we get
@@
where we take the inverse modulo $p^{m-r}$. Hence when $m-r$ is even, then using Proposition 1 and (5.4), we get
@@
Now we consider the case when $m-r$ is odd. We calculate that
$\text{Hesse}f_{l_{1}p^{r},...,l_{n}p^{r},lp^{r}}(y_{1},...,y_{n})=2p^{r}lA.$
@@
$E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})=\left(\frac{\text{det}(2lA)}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{y}}))G_{p}(1)^{n}.$
Similarly as in (5.2), we get
@@
Observe that
$\sum_{\begin{subarray}{c}x=1\\ x\not\equiv 0\bmod p\end{subarray}}^{p^{k}}e_{p^{k}}(xA)=\begin{cases}0,~{}&p^{k-1}\nmid A;\\ -p^{k-1},~{}&p^{k-1}||A;\\ p^{k}-p^{k-1},~{}&p^{k}|A.\end{cases}$
Combining the above equality with (5.7), we get
@@
whenever $p^{m-r-2}\nmid Q^{\prime}(L)$. Otherwise,
$\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})\leq p^{(m+r)n/2+(m-r)}.$

### v5

@@ 3. Proof of Proposition 1
Next, we consider the case when mrm-r is odd, i.e. 2|(mr1)2|(m-r-1). The calculations are similar as above, and writing ${\bf{x}}=p^{(m-r+1)/2}{\bf{z}}+{\bf y}$ with ${\bf{z}}\in\left(\mathbb{Z}_{p}/p^{(m+r-1)/2}\mathbb{Z}_{p}\right)^{n}$ and ${\bf y}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}$, we now obtain the following chain of equations.
$S_{\bf{\alpha}}(f,p^{m})$ $=\sum_{\begin{subarray}{c}{\bf{x}}\in(\mathbb{Z}_{p}/p^{m}\mathbb{Z}_{p})^{n}\\ {\bf{x}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{x}}))$
$=\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m+r-1)/2}\mathbb{Z}_{p})^{n}}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({p^{(m-r+1)/2}\bf{z}}+{\bf y}))$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m+r-1)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(f({\bf{y}})+p^{(m-r+1)/2}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{nr}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m-r-1)/2}\mathbb{Z}_{p})^{n}}e_{p^{(m-r-1)/2}}(p^{-r}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r-1)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})).$
@@ 4. Proof of Theorem 1
$S_{\bf{\alpha}}(f,p^{m})=p^{n(m+r)/2}e_{p^{m}}(f({\bf{\alpha^{*}}})).$
Similarly, if mrm-r is odd, then from (1.1), we obtain
$S_{\bf{\alpha}}(f,p^{m})$ $=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p^{m}}(f({\bf{\alpha}^{*}}+p^{(m-r-1)/2}{\bf{u}}))$
$=p^{(m+r-1)n/2}\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p^{m}}(f({\bf{\alpha}^{*}})+p^{(m-r-1)/2}{\bf{u}}\cdot\nabla f({\bf{\alpha}^{*}})+p^{m-r-1}\overline{2}{\bf{u}}^{t}\text{Hesse}(f)({\bf{\alpha}^{*}}){\bf{u}})$
$=p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(\overline{2}p^{-r}{\bf{u}}^{t}\text{Hesse}(f)({\bf{\alpha}^{*}}){\bf{u}}).$
Take $A=\overline{2}p^{-r}\text{Hesse}(f)({\bf{\alpha}^{*}})$, which is a symmetric matrix. Hence there exists an orthogonal matrix over $\mathbb{F}_{p}$, say $P,$ such that
$PAP^{t}=D=\begin{bmatrix}\lambda_{1}&0&...&0\\ 0&\lambda_{2}&...&0\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&...&\lambda_{n}\end{bmatrix}$
over $\mathbb{F}_{p}$, where $\lambda_{1},...,\lambda_{n}$ are the eigenvalues of $A$ over $\mathbb{F}_{p}$. Let
@@
where $R_{i}$ are the row vectors of $P$. Now ${\bf u}^{t}A{\bf u}={\bf{u}}^{t}P^{t}DP{\bf{u}}=\lambda_{1}(R_{1}{\bf{u}})^{2}+\lambda_{2}(R_{2}{\bf{u}})^{2}+\cdots+\lambda_{n}(R_{n}{\bf{u}})^{2}$. As $P$ is invertible, the map ${\bf{u}}\mapsto P{\bf u}$ is a bijection on $(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}$. Hence we get
$\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(\overline{2}p^{-r}{\bf{u}}^{t}\text{Hesse}(f)({\bf{\alpha}^{*}}){\bf{u}})$ $=\sum_{\begin{subarray}{c}{\bf{u}}\in(\mathbb{Z}_{p}/p\mathbb{Z}_{p})^{n}\end{subarray}}e_{p}(\lambda_{1}(R_{1}{\bf{u}})^{2}+\lambda_{2}(R_{2}{\bf{u}})^{2}+\cdots+\lambda_{n}(R_{n}{\bf{u}})^{2})$
@@
$S_{\bf{\alpha}}(f,p^{m})=\left(\frac{\text{det}(\overline{2}p^{-r}\text{Hesse}(f)({\bf{\alpha}^{*}}))}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))G_{p}(1)^{n},\quad\mbox{if }m-r\mbox{ is odd,}$
@@ 5.1. Multiple Poisson summation
Setting $x_{i}=y_{i}+k_{i}p^{m}$, where $k_{i}\in\mathbb{Z}$ and $1\leq y_{i}\leq p^{m}$, we apply Proposition 2 to obtain
@@ 5.2. Evaluation of exponential sums
(5.3) $f_{k_{1},...,k_{n},h}(y_{1},y_{2},...,y_{n}):=h\sum_{i,j=1}^{n}a_{ij}y_{i}y_{j}+\sum_{i=1}^{n}k_{i}y_{i}.$
The contributions of r=m1r=m-1 and $r=m$ to the right-hand side of (5.2) are $O_{\varepsilon}(1)$ if $N\geq p^{m\varepsilon}$ by the rapid decay of $\hat{\Phi}$ as $(k_{1},...,k_{n})\neq(0,...,0)$. We fix $l_{i}=k_{i}/p^{r}$. From (5.3) we calculate
$\nabla f_{k_{1},...,k_{n},h}=\left(h\frac{\partial Q}{\partial y_{1}}+p^{r}l_{1},...,h\frac{\partial Q}{\partial y_{n}}+p^{r}l_{n}\right).$
In the following, we write $f=f_{k_{1},...,k_{n},h}$ for simplicity. If $\text{\rm ord}_{p}(h)>r$ then $\text{\rm ord}_{p}(\nabla(f))=r$. So $p^{-r}\nabla(f)\equiv(l_{1},...l_{n})\bmod p,$ where $(l_{1},...,l_{n})\not\equiv(0,...,0)\bmod p$. Hence there is no critical point of $f_{k_{1},...,k_{n},h}$. Now using $(i)$ of Theorem 1, we have $E(k_{1},...,k_{n},h;p^{m})=0$. On the other hand if $\text{\rm ord}_{p}(h)<r$ then $\text{\rm ord}_{p}(\nabla(f))=\text{\rm ord}_{p}(h)$. Hence $p^{-\text{\rm ord}_{p}(h)}\nabla(f)\equiv{\boldsymbol{0}}\bmod p,$ implies $AY\equiv{\boldsymbol{0}}\bmod p$ for $Y=(y_{1}~y_{2}~\cdots~y_{n})^{t}$. Since $A$ is non-singular, Y=(0,,0)Y=(0,...,0) is the only critical point in this case. Hence $E(k_{1},...,k_{n},h;p^{m})=0.$ Thus from now onward we consider $\text{\rm ord}_{p}(h)=r$. Therefore $p^{\text{\rm ord}_{p}(\nabla(f))}=p^{r}=(h,p^{m})$ and we write $l=h/p^{r}$. So if $N\geq p^{m\varepsilon}$, we have
(5.4) $U=\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)e_{p^{m}}\left(-\sum_{i=1}^{n}l_{i}p^{r}x_{0,i}\right)\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})+O_{\varepsilon}(1).$
@@
(5.5) $\frac{\partial f_{l_{1}p^{r},...,l_{n}p^{r},lp^{r}}}{\partial y_{i}}(y_{1},y_{2},...,y_{n})=p^{r}\left(l\frac{\partial{Q}}{\partial{y_{i}}}+l_{i}\right),~\forall~i.$
In the following, we will apply Proposition 1 to evaluate the exponential sums $E(l_{1}p^{r},...,l_{n}p^{r};p^{m})$. Now consider the case when mrm-r is even. Note that $A$ is the symmetric matrix associated with the quadratic form $Q(y_{1},...,y_{n})=\sum_{i,,j=1}^{n}a_{ij}y_{i}y_{j}$. Hence $\frac{\partial Q}{\partial y_{j}}=2(a_{j1}y_{1}+a_{j2}y_{2}+\cdots+a_{jn}y_{n})=2(a_{j1}~a_{j2}~\cdots~a_{jn})Y$. Therefore $p^{-r}\nabla f_{k_{1},...,k_{n},h}(y_{1},...,y_{n})\equiv\boldsymbol{0}\bmod p^{m-r}$ implies
$2lAY+L\equiv\boldsymbol{0}\bmod p^{m-r}.$
As $A$ is non-singular modulo $p$, we have
@@
where $Y=(y_{1}~y_{2}~\cdots~y_{n})^{t}$ and $L=(l_{1}~l_{2}~\cdots~l_{n})^{t}$. This is the only critical point for $f_{k_{1},...,k_{n},h}(y_{1},y_{2},...,y_{n}).$ Now putting the values of $y_{i}$ in (5.3), we get
@@
where we take the inverse modulo $p^{m-r}$. Hence when mrm-r is even, then using Proposition 1, we get
@@
Now we consider the case when mrm-r is odd. We calculate that
$\text{Hesse}(f_{l_{1}p^{r},...,l_{n}p^{r},lp^{r}})(y_{1},...,y_{n})=2p^{r}lA.$
@@
$E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})=\left(\frac{\text{det}(2lA)}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{y}}))G_{p}(1)^{n},$
where ${\bf y}=Y^{t}$. Similarly as in (5.6), we get
@@
Observe that for any $a\in\mathbb{Z}$, we have
$\sum_{\begin{subarray}{c}x=1\\ x\not\equiv 0\bmod p\end{subarray}}^{p^{k}}e_{p^{k}}(xa)=0\quad\mbox{ if }p^{k-1}\nmid a$
and
$\sum_{\begin{subarray}{c}x=1\\ x\not\equiv 0\bmod p\end{subarray}}^{p^{k}}\left(\frac{x}{p}\right)e_{p^{k}}(xa)$ $=\sum\limits_{b=1}^{p-1}\left(\frac{b}{p}\right)\sum_{\begin{subarray}{c}x=1\\ x\equiv b\bmod p\end{subarray}}^{p^{k}}e_{p^{k}}(xa)$
$=\sum\limits_{b=1}^{p-1}\left(\frac{b}{p}\right)\sum_{\begin{subarray}{c}y=1\end{subarray}}^{p^{k-1}}e_{p^{k}}((py+b)a)$
$=\sum\limits_{b=1}^{p-1}\left(\frac{b}{p}\right)e_{p^{k}}(ba)\sum_{\begin{subarray}{c}y=1\end{subarray}}^{p^{k-1}}e_{p^{k-1}}(ya)$
$=0\quad\mbox{ if }p^{k-1}\nmid a.$
Using the above equalities, we deduce from (5.7) that
@@
whenever $p^{m-r-1}\nmid Q^{\prime}(L)$. Otherwise, a trivial estimate gives
$\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}|E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})|\leq p^{(m+r)n/2+(m-r)}.$